Find the principle values of the natural logarithm of z =(1+1)* Express in the form a tib
Answer #1


Given atib= (1+) arib=(1+3)*(1+3) a tib= (1 +5² +22) (14.3 + 2) 3-1 that we know atib= (-1 +23) (1-1 +23) atib = (25) (23) at